OpenStudy (anonymous):

solve for X... (2x-8)(x+10)=0

OpenStudy (anonymous):

x = 4 or -10

OpenStudy (anonymous):

how about for.... (-4x+16)(5x+10)=0

OpenStudy (anonymous):

x=4,-1/5

OpenStudy (anonymous):

(2x-8)(x+10)=0 2x-8=0 x=4 x+10=0 x=-10

OpenStudy (anonymous):

(-4x+16)(5x+10)=0 -4x+16=0 x=4 5x+10=0 x=-2

OpenStudy (anonymous):

thnx u guys :)